OpenStudy (anonymous):

|4x - 10| - 16 = -12

OpenStudy (anonymous):

x= 1.5,3.5

OpenStudy (anonymous):

Move the 16 to the other side |4x-10|=4 So, 4x-10=4 or 4x-10=-4. Solve both for x.